# Preferences Evolve And So Should Your Bandits: Bandits with Evolving States for Online Platforms

### v2

@@ 12      Call Algorithm 2 for $\lambda=\hat{\lambda}_{1}$
We first prove prove $|q_{t}\left(H_{1:t-1}\right)-q_{t}\left(H_{1:t-1}^{\prime}\right)|\leq\sqrt{\log T}/(T/K)^{1/3}$. If $\max_{i,j\in[K]}\{|b_{i}-b_{j}|\}\leq\sqrt{\log T}/(T/K)^{1/3}$ then for all $i,j$ it applies $|b_{i}-b_{j}|\leq\sqrt{\log T}/(T/K)^{1/3}$.
$|q_{t}\left(H_{1:t-1}\right)-q_{t}\left(H_{1:t-1}^{\prime}\right)|$ $=(1-\lambda)^{t}+\lambda\cdot\sum_{s=0}^{t-1}(1-\lambda)^{t-1-s}\cdot b_{I_{s}}-(1-\lambda)^{t}+\lambda\cdot\sum_{s=0}^{t-1}(1-\lambda)^{t-1-s}\cdot b_{I_{s}}^{\prime}$
$=\lambda\sum_{s=0}^{t-1}(1-\lambda)^{t-1-s}\cdot b_{I_{s}}-\lambda\sum_{s=0}^{t-1}(1-\lambda)^{t-1-s}\cdot b_{I_{s}}^{\prime}=\lambda\sum_{s=0}^{t-1}(1-\lambda)^{t-1-s}\cdot(b_{I_{s}}-b_{I_{s}}^{\prime})$
$\leq\left(1-(1-\lambda)^{t}\right)(b_{I_{s}}-b_{I_{s}}^{\prime})\leq\sqrt{\log T}/(T/K)^{1/3}.{}$ (31)
Since, $|b_{i}-b_{j}|\leq\sqrt{\log T}/(T/K)^{1/3}$ for all $i,j$ and  31 we get $|Y_{i}-Y_{i,j}|\leq\frac{\sqrt{\log T}}{(T/K)^{1/3}}$ and so on $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|\leq\frac{3\sqrt{\log T}}{(T/K)^{1/3}}$. Thus, the algorithm will call $EXP3.P$ and the regret will be:
$R_{\textsc{DES}}(T)$ $=\mathbb{E}\left[\sum_{t\in[T]}r_{\pi_{t}^{\star}}\cdot q_{t}\left(H_{1:t-1}^{\pi^{\star}}\right)-\sum_{t\in[T]}r_{I_{t}}\cdot q_{t}\left(H_{1:t-1}^{\textsf{ALG}}\right)\right]+2K^{1}/3T^{2/3}$
$=\mathbb{E}\left[\sum_{t\in[T]}r_{\pi_{t}^{\star}}\cdot q_{t}\left(H_{1:t-1}^{\pi^{\star}}\right)-r_{i^{\star}}q_{t}\left(H_{1:t-1}^{i^{\star}}\right)-\sum_{t\in[T]}r_{I_{t}}\cdot q_{t}\left(H_{1:t-1}^{\textsf{ALG}}\right)-r_{i^{\star}}q_{t}\left(H_{1:t-1}^{i^{\star}}\right)\right]+2K^{1/3}T^{2/3}$
$\leq\sum_{t\in[T]}r_{\pi_{t}^{\star}}\cdot q_{t}\left(H_{1:t-1}^{\pi^{\star}}\right)-r_{i^{\star}}q_{t}\left(H_{1:t-1}^{i^{\star}}\right)+\mathcal{O}(\sqrt{KT\log T})+2K^{1/3}T^{2/3}$
$\leq\sum_{t\in[T]}r_{i^{\star}}\left(\cdot q_{t}\left(H_{1:t-1}^{\pi^{\star}}\right)-q_{t}\left(H_{1:t-1}^{i^{\star}}\right)\right)+\mathcal{O}(\sqrt{KT\log T})+2K^{1/3}T^{2/3}$
$\leq\sum_{t\in[T]}r_{i^{\star}}\frac{K^{1/3}}{T^{1/3}}+\mathcal{O}(\sqrt{KT\log T})+2K^{1/3}T^{2/3}\leq\mathcal{O}(K^{1/3}T^{2/3})$
where $r_{i^{\star}}$ is the $r_{i}$ of the best fixed arm. ∎
Theorem E.9.
If $\lambda>O(K^{1/3}/T^{1/3})$ and $\max_{i,j\in[K]}\{r_{i}(b_{i}-b_{j})\}>\omega(\sqrt{\log T}/(T/K)^{2/3})$ then Algorithm 8 calls Algorithm 7 and achieves regret $R_{\textsc{DES}}(T)=\mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right)$
@@
First we prove that exist $i,j$ $|Y_{i}-Y_{i,j}|\geq\Omega(\sqrt{\log T}K^{1/3}/T^{1/3})$ and so the algorithm will not run $EXP3.P$. If $\lambda>K^{1/3}/T^{1/3}$ the $1/\lambda T^{2/3}<K^{1/3}/T^{1/3}$ and so on for all arms $i,j$ in $[K]$:
$\left|(1-\lambda)^{(T/K)^{2/3}}\frac{1-(1-\lambda)^{(T/K)^{2/3}}(q_{s,i}-b_{i})}{\lambda T^{2/3}}\right|<\frac{K^{1/3}}{T^{1/3}}$
and
$\left|\frac{(1-\lambda)^{2(T/K)^{2/3}}}{(T/K)^{2/3}}\left(r_{i}q_{s,(i,j)}-\frac{r_{i}b_{j}}{2-\lambda}-(1-\lambda)\frac{r_{i}b_{i}}{2-\lambda}\right)\frac{1-(1-\lambda)^{2(T/K)^{2/3}}}{1-(1-\lambda)^{2}}\right|<\frac{K^{1/3}}{T^{1/3}}$
Thus, $Y_{i}-Y_{(i,j)}\geq\left|r_{i}b_{i}-r_{i}\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}\right|-2\frac{K^{1/3}}{T^{1/3}}$. Let $i^{\star},j^{\star}$ be two arms that satisfy $(r_{i^{\star}}b_{i^{\star}}-b_{j^{\star}})\geq\frac{\sqrt{\log T}K^{1/3}}{T^{1/3}}$ then for the random arm $j$ that we choose in line $5$ of Algorithm 8 applies $\max\{|b_{i^{\star}}-b_{j}|,|b_{j^{\star}}-b_{j}|\}\geq O(\sqrt{\log T}/(T/K)^{2/3})/2=O(\sqrt{\log T}/(T/K)^{2/3})$ Thus, for that pair:
$|Y_{i}-Y_{i,j}|$ $\geq r_{i}\left|b_{i}-\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}\right|-\frac{2}{T^{1/3}}$
$=r_{i}\left|\frac{b_{i}-b_{j}}{2-\lambda}\right|-\frac{2}{T^{1/3}}\geq\Omega(\sqrt{\log T}/(T/K)^{1/3})$
Thus, the Algorithm 8 goes to else. Then, we call Algorithm 7 for $\tilde{N}(\lambda)=\log T$ and for $\tilde{N}(\lambda)=2\log T$ and we keep calling it by doubling the $N(\lambda)$s until $|\hat{\lambda}_{1}-\hat{\lambda}_{2}|\leq\delta$. Assume, $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ be the $\lambda$’s when event $|\hat{\lambda}_{1}-\hat{\lambda}_{2}|\leq\delta$ occurs. Now we use lemma E.4 to bound $|\hat{\lambda}_{1}-\hat{\lambda}_{2}|$, also observe that the assumptions we made in the statement of this lemma applies as $\lambda>O(K^{1/3}/T^{1/3})$.(For ease of the analysis assume $A=r_{i}(1-\lambda)^{\tilde{N}_{1}(\lambda)}(q_{0,i}-b_{i})+r_{i}b_{i}-r_{i}(1-\lambda)^{\tilde{N}_{1}(\lambda)}(q_{0,j}-b_{i,j})-r_{i}\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}$ and $B=r_{i}(1-\lambda)^{\tilde{N}_{2}(\lambda)}(q_{0,i}-b_{i})+r_{i}b_{i}-r_{i}(1-\lambda)^{\tilde{N}_{2}(\lambda)}(q_{0,j}-b_{i,j})-r_{i}\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}$, where $b_{i,j}=\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}$.)
$|\hat{\lambda}_{1}-\hat{\lambda}_{2}|$ $=\left|1-\frac{\hat{r}_{i,b_{j}}^{1}-\hat{r}_{i,j}^{1}}{\hat{r}_{i,b_{i}}^{1}-\hat{r}_{i,j}^{1}}-1+\frac{\hat{r}_{i,b_{j}}^{2}+\hat{r}_{i,j}^{2}}{\hat{r}_{i,b_{i}}^{2}-\hat{r}_{i,j}^{2}}\right|$
$\geq\left|\frac{\mathbb{E}[\hat{r}_{i,b_{j}}^{1}]-\mathbb{E}[\hat{r}_{i,j}^{1}]}{\mathbb{E}[\hat{r}_{i,b_{i}}^{1}]-\mathbb{E}[\hat{r}_{i,j}^{1}]}-\frac{\mathbb{E}[\hat{r}_{i,b_{j}}^{2}]-\mathbb{E}[\hat{r}_{i,j}^{2}]}{\mathbb{E}[\hat{r}_{i,b_{i}}^{2}]-\mathbb{E}[\hat{r}_{i,j}^{2}]}\right|-\Omega(\delta)$
$\geq\bigg{|}\frac{r_{i}(1-\lambda)^{\tilde{N}_{1}(\lambda)}(q_{0,b_{j}}-b_{j})+r_{i}b_{j}-r_{i}(1-\lambda)^{\tilde{N}_{1}(\lambda)}(q_{0,j}-b_{i,j})-r_{i}\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}}{r_{i}(1-\lambda)^{\tilde{N}_{1}(\lambda)}(q_{0,i}-b_{i})+r_{i}b_{i}-r_{i}(1-\lambda)^{\tilde{N}_{1}(\lambda)}(q_{0,j}-b_{i,j})-r_{i}\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}}$
$-\frac{r_{i}(1-\lambda)^{\tilde{N}_{2}(\lambda)}(q_{0,b_{j}}-b_{j})+r_{i}b_{j}-r_{i}(1-\lambda)^{\tilde{N}_{2}(\lambda)}(q_{0,j}-b_{i,j})-r_{i}\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}}{r_{i}(1-\lambda)^{\tilde{N}_{2}(\lambda)}(q_{0,i}-b_{i})+r_{i}b_{i}-r_{i}(1-\lambda)^{\tilde{N}_{2}(\lambda)}(q_{0,j}-b_{i,j})-r_{i}\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}}\bigg{|}-\Omega(\delta)$
$\geq\bigg{|}\frac{r_{i}(1-\lambda)^{\tilde{N}_{1}(\lambda)}(q_{0,b_{j}}-b_{j})-r_{i}(1-\lambda)^{\tilde{N}_{1}(\lambda)}(q_{0,j}-b_{i,j})}{\max(A,B)}$
$-\frac{r_{i}(1-\lambda)^{\tilde{2}N_{1}(\lambda)}(q_{0,b_{j}}-b_{j})-r_{i}(1-\lambda)^{2\tilde{N}_{1}(\lambda)}(q_{0,j}-b_{i,j})}{\max\{A,B\}}\bigg{|}$
$\geq\left|\frac{(1-\lambda)^{\tilde{N}_{1}(\lambda)}\left(r_{i}(q_{0,b_{j}}-b_{j})\left(1-(1-\lambda)^{\tilde{N}_{1}(\lambda)}\right)-r_{i}(q_{0,j}-b_{i,j})\left(1-(1-\lambda)^{\tilde{N}_{1}(\lambda)}\right)\right)}{\max\{A,B\}}\right|-\Omega(\delta)$
Because $\lambda>\frac{K^{1/3}}{T^{1/3}}$ solving the equation we get $(1-\lambda)^{N_{1}(\lambda)}\leq|\Omega(\delta)|$ and also that means that:
$\hat{\lambda}_{1}$ $\leq 1-\frac{|\mathbb{E}[\hat{r}_{i,b_{j}}]-\mathbb{E}[\hat{r}_{i,j}]|-2\delta}{|\mathbb{E}[\hat{r}_{i,b_{i}}]-\mathbb{E}[\hat{r}_{i,j}]|+2\delta}$
$=1+\frac{r_{i}b_{j}-r_{i}\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}}{r_{i}b_{i}-r_{i}\frac{b_{j}+(1-\lambda)b_{i}}{2-\lambda}}+O(\delta)$
$=1-1+\lambda+O(\delta)$
$=\lambda+O(\delta)$
and $\hat{\lambda}_{1}\geq\lambda-\Omega(\delta)$ respectively. Then we call Algorithm 2 for $\lambda=\hat{\lambda}_{1}$ the Theorem 3.1 applies and we get an additional $O(\delta)T$ from $\hat{\lambda}_{1}$. From the exploration of Algorithm 8 we get an additional $K^{1/3}T^{2/3}$ regret and from Algorithm 7 an $N(\lambda)\log T$. Thus, the total regret is:
@@
$R_{\textsc{DES}}(T)$ $=\mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right)+2K^{1/3}T^{2/3}+O(\delta)\cdot T+N(\lambda)\log T$
@@
$=\mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right)$

### v4

@@ E.2.2 Analysis
First, we prove that if the assumptions of the lemma hold (i.e., $\max_{i,j\in[K]}\{r_{i}\cdot|b_{i}-b_{j}|\}>\omega(\sqrt{\log T}/(T/K)^{1/3})$ and $\lambda>\omega((K/T)^{2/3})$), then the difference between estimators $\hat{\mu}_{i}$ and $\hat{\mu}_{i,j}$ is sufficiently large, i.e., $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|>5(\sqrt{\log T}/(T/K)^{1/3})$ with probability $1-\frac{2}{T}$. In this case, Algorithm 8 will not call EXP3.P. From lemmas E.4 E.5, we get that with probability $1-2/T$:
$\hat{\mu}_{i}-\hat{\mu}_{i,j}\leq 2\frac{\sqrt{\log T}}{(T/K)^{1/3}}+\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$
which can be written also as:
$\hat{\mu}_{i,j}-\hat{\mu}_{i}\geq-2\frac{\sqrt{\log T}}{(T/K)^{1/3}}+\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]$ (35)
and also:
$\hat{\mu}_{i}-\hat{\mu}_{i,j}\geq-2\frac{\sqrt{\log T}}{(T/K)^{1/3}}+\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\frac{1}{(T/K)^{2/3}}\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right].$ (36)
So, we need to bound the difference $\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$:
$\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]$ $=\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot\left((1-\lambda)^{(T/K)^{2/3}+s}\cdot(q_{0,i}-b_{i})+b_{i}\right)$
$\quad\quad-\sum_{s=0}^{(T/K)^{2/3}-1}r_{i}\cdot(1-\lambda)^{(T/K)^{2/3}+2s}\left(q_{0,i}^{\text{ALT}}-W_{i,j}(\lambda)\right)+r_{i}W_{i,j}(\lambda)$
$=\sum_{s=0}^{(T/K)^{2/3}-1}\underbrace{r_{i}\left((1-\lambda)^{(T/K)^{2/3}+s}q_{0,i}-(1-\lambda)^{(T/K)^{2/3}+2s}q_{0,i}^{ALT}\right)}_{A_{1}}$
$+\underbrace{r_{i}b_{i}-r_{i}W_{i,j}(\lambda)}_{A_{2}}+\underbrace{r_{i}W_{i,j}(\lambda)(1-\lambda)^{(T/K)^{2/3}+2s}-r_{i}b_{i}(1-\lambda)^{(T/K)^{2/3}+s}}_{A_{3}}$
The quantity $(1-\lambda)^{T^{2/3}}\leq\frac{1}{T}$, thus $A_{1}\leq T^{1/3}$ and $A_{3}\leq T^{1/3}$. Now fix a pair $(i,j)\in[K]^{2}$ chosen such that: $j$ is uniformly-at-random chosen among the $K$ arms and $i$ is chosen such that $i=\arg\max_{i\in[K]}r_{i}\cdot|b_{i}-b_{j}|$. Then, our assumption $r_{i}\cdot|b_{i}-b_{j}|>\omega(\sqrt{\log T}/(T/K)^{1/3})$ (since this is the sub-case that we are currently investigating) implies that:
$r_{i}|b_{i}-W_{i,j}|=\left|r_{i}b_{i}-r_{i}b_{i}\frac{1-\lambda}{2-\lambda}-r_{i}b_{j}\frac{1}{2-\lambda}\right|$ $=\left|\frac{r_{i}b_{i}-r_{i}b_{j}}{2-\lambda}\right|$
$\geq|r_{i}b_{i}-r_{i}b_{j}|/2>\omega(\sqrt{\log T}/(T/K)^{1/3}).{}$ (37)
Thus, $|\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}\right]-\mathbb{E}\left[\sum_{s=0}^{(T/K)^{2/3}-1}R_{s,i}^{\text{ALT}}\right]|>\omega\left(\sqrt{\log T}(T/K)^{1/3}\right)$. Putting this result into equations 35 and 36 we get that $|\hat{\mu}_{i}-\hat{\mu}_{i,j}|>\omega\left(\sqrt{\log T}/(T/K)^{1/3}\right)>5\sqrt{\log T}/(T/K)^{1/3}$ with probability at least $1-2/T$.
Hence, the if clause of Algorithm 8 in line $4$ is not satisfied, and we move to line $6$ with probability $1-2/T$. Now we are going to prove that the while loop of Algorithm 8 in line $9$ breaks when $\tilde{N}\geq N(\lambda)$ and so on the Algorithm 7 is called and also that for $\tilde{N}\leq\frac{N(\lambda)}{\log T}$ the while loop never breaks, with high probability. Let’s start with the second; we prove that if $\tilde{N}\leq\frac{N(\lambda)}{\log T}$ then the while loop never breaks with high probability using contradiction. The algorithm at each loop builds an estimator $\tilde{\mu}_{i}$, for which using Hoeffding inequality we get that:
$\Pr\left[\left|\tilde{\mu}_{i}-\tilde{Y}_{i}\right|\geq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\leq\frac{1}{T^{2}},$ (38)
where $\tilde{Y}_{i}=r_{i}\tilde{q}_{s,i}=r_{i}(1-\lambda)^{\tilde{N}}(q_{j}-b_{i})+r_{i}b_{i}$, $q_{s,i}$ is the value of the state before starting sampling in line $14$, and $q_{j}$ is the value of the state after playing arm $j$ for $\tilde{N}$ rounds. Taking a union bound on all loops (there are at most $\log T$ of them), we get:
$\Pr\left[\left|\tilde{\mu}_{i}-\tilde{Y}_{i}\right|\geq\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right]\leq\frac{\log T}{T^{2}}.$
Let $\hat{N}$ be the value of $\tilde{N}$ when the while loop breaks. Now we bound the difference of the estimators $|\hat{\mu}_{i}-\tilde{\mu}_{i}|$ to find the $\hat{N}$:
$|\hat{\mu}_{i}-\tilde{\mu}_{i}|$ $=\left|\hat{\mu}_{i}-r_{i}b_{i}+r_{i}b_{i}-\tilde{\mu}_{i}+\tilde{Y}_{i}-\tilde{Y}_{i}\right|$
$=\left|\hat{\mu}_{i}-r_{i}b_{i}+r_{i}b_{i}-\tilde{\mu}_{i}+\tilde{Y}_{i}-r_{i}(1-\lambda)^{\hat{N}}(q_{j}-b_{i})-r_{i}b_{i}\right|$
$=\left|(\hat{\mu}_{i}-r_{i}b_{i})+(\tilde{\mu}_{i}-\tilde{Y}_{i})-r_{i}(1-\lambda)^{\hat{N}}(q_{j}-b_{i})\right|{}$ (39)
For $\hat{N}\leq N(\lambda)/\log T$, the quantity $(1-\lambda)^{\hat{N}}\geq(1-\lambda)^{1/\log(1-\lambda)}$, which is constant. Also remember that $r_{i}\cdot|q_{j}-b_{i}|=\Theta(r_{i}|b_{j}-b_{i}|)>\omega(\sqrt{\log T}/(T/K)^{1/3})$ (from assumption), $|\hat{\mu}_{i}-r_{i}b_{i}|\leq\sqrt{\log T}/(K/T)^{1/3}$ and $|\tilde{\mu}_{i}-\tilde{Y}_{i}|\leq\sqrt{\log T}/(K/T)^{1/3}$, with probability $(1+\log T)/T^{2}$. Putting all together in 39 we get:
$|\hat{\mu}_{i}-\tilde{\mu}_{i}|>\omega\left(\frac{\sqrt{\log T}}{(T/K)^{1/3}}\right)>3\sqrt{\log T}/(T/K)^{1/3}$
Now we prove that: $|\hat{\mu}_{i}-\tilde{\mu}_{i}|\leq 3\sqrt{\log T}/(T/K)^{1/3}$ if $\hat{N}=N(\lambda)$. For $\hat{N}=\log(T)/\log(1-\lambda)$, equation 39 becomes:
$|\hat{\mu}_{i}-\tilde{\mu}_{i}|$ $=\left|(\hat{\mu}_{i}-r_{i}b_{i})+(\tilde{\mu}_{i}-\tilde{Y}_{i})-r_{i}(1-\lambda)^{\frac{\log(T)}{\log(1-\lambda)}}(q_{j}-b_{i})\right|$
$\leq\left|(\hat{\mu}_{i}-r_{i}b_{i})\right|+\left|(\tilde{\mu}_{i}-\tilde{Y}_{i})\right|+\left|r_{i}(1-\lambda)^{\frac{\log(T)}{\log(1-\lambda)}}(q_{j}-b_{i})\right|$
$\leq 2\frac{\sqrt{\log T}}{(T/K)^{1/3}}+\frac{1}{T}\leq 3\sqrt{\log T}/(T/K)^{1/3}$ (with probability $(\log T+1)/T^{2}$)
Thus, the $while$ breaks when the $\tilde{N}$ overcomes the value of $N(\lambda)$. Algorithm keeps doubling the value of $\tilde{N}$, hence with probability $(\log T+1)/T^{2}$, $\hat{N}(\lambda)\leq 2N(\lambda)$. Also $\hat{N}\geq N(\lambda)/\log T$ with probability $(\log T+1)/T^{2}$. So, combining these two by using union bound we have that Algorithm 7 is called with $\frac{N(\lambda)}{\log T}\leq\tilde{N}\leq 2N(\lambda)$ with probability $1-2(\log T+1)/T^{2}$. ∎
Lemma E.9.
If $\tilde{N}\geq N(\lambda)$, then the estimator $\hat{\lambda}$ that Algorithm 7 builds is such that:
$-4\delta-\varepsilon\leq\left|\hat{\lambda}-\lambda\right|\leq 4\delta+\varepsilon,$
with probability at least $1-6\exp(-2M\delta^{2})$, for $\varepsilon,M$ as defined in Theorem 3.1.
Proof.
From Lemma 3.6, the values of the states $q_{i},q_{j}$ and $q_{i,j}^{\text{ALT}}$ after playing $\tilde{N}\geq N(\lambda)$ rounds repeatedly arm $i$, arm $j$ and alternately arms $i,j$ respectively in Algorithm 7 in lines $3$, $6$ and $11-12$ becomes:
$|q_{i}-b_{i}|\leq\varepsilon\quad\quad|q_{j}-b_{j}|\leq\varepsilon\quad\quad and\quad\quad\left|q_{i,j}^{\text{ALT}}-W_{i,j}(\lambda)\right|\leq\varepsilon.$
The way the estimators are built in Algorithm 7 is the same as in Algorithm 2. Thus, we adapt the results of Lemma 3.9 on these estimators to avoid repeating the same proofs:
$\Pr\left[\left|\hat{r}_{i,i}-r_{i}b_{i}\right|\geq\delta\right]\leq 2\exp(-2M\delta^{2})$
$\Pr\left[\left|\hat{r}_{i,j}-r_{i}b_{j}\right|\geq\delta\right]\leq 2\exp(-2M\delta^{2})$
$\Pr\left[\left|\hat{r}_{i,j}^{\text{ALT}}-r_{i}b_{i}\frac{1-\lambda}{2-\lambda}-r_{i}b_{j}\frac{1}{2-\lambda}\right|\geq\delta\right]\leq 2\exp(-2M\delta^{2})$
Taking a union bound, we get that all three inequalities hold with probability at most $6\exp(-2M\delta^{2})$. For the rest of the analysis, we condition on the event that all three inequalities hold. Thus, we have a system of equations and solving for $\hat{\lambda}$ we get:
$\hat{\lambda}$ $=\frac{\hat{r}_{i,i}+\hat{r}_{i,j}-2\hat{r}_{i,j}^{\text{ALT}}}{\hat{r}_{i,i}-\hat{r}_{i,j}^{\text{ALT}}}$
$\leq 1+\frac{\hat{r}_{i,j}-\hat{r}_{i,j}^{\text{ALT}}}{\hat{r}_{i,i}-\hat{r}_{i,j}^{\text{ALT}}}\leq 1+\frac{r_{i}b_{j}-r_{i}b_{i}\frac{1-\lambda}{2-\lambda}-r_{i}b_{j}\frac{1}{2-\lambda}}{r_{i}b_{i}-r_{i}b_{i}\frac{1-\lambda}{2-\lambda}-r_{i}b_{j}\frac{1}{2-\lambda}}+4\delta$ ($r_{i}|b_{i}-b_{j}|>>\delta$)
$=1+\frac{(2-\lambda)r_{i}b_{j}-r_{i}b_{i}(1-\lambda)-r_{i}b_{j}}{r_{i}b_{i}(2-\lambda)-r_{i}b_{i}(1-\lambda)-r_{i}b_{j}}+4\delta$
$=1+\frac{(r_{i}b_{j}-r_{i}b_{i})(1-\lambda)}{r_{i}b_{i}-r_{i}b_{j}}+4\delta=\lambda+4\delta.$
We do the same same to lower bound the $\hat{\lambda}$:
$\hat{\lambda}$ $=\frac{\hat{r}_{i,i}+\hat{r}_{i,j}-2\hat{r}_{i,j}^{\text{ALT}}}{\hat{r}_{i,i}-\hat{r}_{i,j}^{\text{ALT}}}$
$\geq 1+\frac{\hat{r}_{i,j}-\hat{r}_{i,j}^{\text{ALT}}}{\hat{r}_{i,i}-\hat{r}_{i,j}^{\text{ALT}}}\geq 1+\frac{r_{i}b_{j}-r_{i}b_{i}\frac{1-\lambda}{2-\lambda}-r_{i}b_{j}\frac{1}{2-\lambda}}{r_{i}b_{i}-r_{i}b_{i}\frac{1-\lambda}{2-\lambda}-r_{i}b_{j}\frac{1}{2-\lambda}}-4\delta$ ($r_{i}|b_{i}-b_{j}|>>\delta$)
$=1+\frac{(2-\lambda)r_{i}b_{j}-r_{i}b_{i}(1-\lambda)-r_{i}b_{j}}{r_{i}b_{i}(2-\lambda)-r_{i}b_{i}(1-\lambda)-r_{i}b_{j}}-4\delta$
$=1+\frac{(r_{i}b_{j}-r_{i}b_{i})(1-\lambda)}{r_{i}b_{i}-r_{i}b_{j}}-4\delta=\lambda-4\delta.$
Thus, we have:
$\Pr\left[\left|\hat{\lambda}-\lambda\right|\geq 4\delta\right]\leq 6\exp(-2M\delta^{2}).$
We are now finally ready to provide the full proof of Theorem E.2.
@@
For the case where $\lambda\in[0,\tilde{\Theta}(1/T)]$ we have already proven the results in Lemma E.6. For the case where $\lambda>\omega((K/T)^{2/3})$ we have two sub-cases. For the sub-case $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}\leq O(\sqrt{\log T}/(T/K)^{1/3})$ from Lemma E.7 it holds that:
@@
$R_{\textsc{DES}}(T)\leq\widetilde{\mathcal{O}}(K^{1/3}T^{2/3})\leq\mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right)$
Now to prove that:
$R_{\textsc{DES}}(T)\leq\mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right),$
for $\lambda>\omega((K/T)^{2/3})$, it remains to prove that Algorithm 8 achieves regret
$R_{\textsc{DES}}(T)\leq\mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right),$
for the sub-case $\lambda>\omega((K/T)^{2/3})$ and $\max_{i,j\in[K]}\{r_{i}|b_{i}-b_{j}|\}>\omega(\sqrt{\log T}/(T/K)^{1/3})$. In that case from Lemma E.8, Algorithm 7 is called, for $\tilde{N}$ such that $\frac{N(\lambda)}{\log T}<\tilde{N}\leq 2N(\lambda)$ with probability at least $1-2\frac{\log T}{T}$. Hence, Algorithm 7 runs for $\tilde{N}=\tilde{N}\cdot\log T>N(\lambda)$ with probability at least $1-2\frac{\log T}{T}$. As a result, Lemma E.9 applies and Algorithm 2 is called for the estimator $\hat{\lambda}$, with $|\lambda-\hat{\lambda}|\leq\delta$ with probability at $1-6\exp(-2M\delta^{2})$, where $M$ is as defined in Theorem 3.1. Using a union bound for the case that both events $|\lambda-\hat{\lambda}|\leq\delta$ and $\frac{N(\lambda)}{\log T}<\tilde{N}\leq 2N(\lambda)$ we have:
  • $K^{1/3}T^{2/3}$ regret for the exploration phase of Algorithm 8 (Lines 1–3),
  • at most $2N(\lambda)T^{2/3}\cdot\log T$ for the binary search part of the Algorithm 8 (Lines 9–15),
  • $3MN(\lambda)$ from Algorithm 7
@@
  • $\mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right)$ from Algorithm 2 (see Theorem 3.1).
Let $A$ be the sum of all regret quantities mentioned above, then the total expected regret (accounting for the probabilities of all the random events occurring) will be equal to:
$R_{\textsc{DES}}(T)$ $\leq A\left(1-2\frac{\log T+1}{T}-6\exp(-2M\delta^{2})\right)+\left(2\frac{\log T+1}{T}+6\exp(-2M\delta^{2})\right)T$
$\leq\mathcal{O}\left(\left(\frac{K\log(T)\log(\lambda)}{\log(1-\lambda)}\right)^{1/3}\cdot T^{2/3}\right)$